OpenStudy (anonymous):

what am i doing wrong here? i thought i understood this. Use the parametric equations of an ellipse x=10cosθ y=8sinθ 0≤θ≤2π to find the area that it encloses. did the dx/dθ and got -10sinθ.

OpenStudy (anonymous):

so i get \[\int\limits_{0}^{2\pi}-10\sin \theta*8\sin \theta \] \[\int\limits_{0}^{2\pi} -80\sin ^{2} \theta \] \[-40(\theta-\frac{ \sin2\theta }{ 2})\] plug in 2pi for theta...but webwork says it's wrong. what did i do wrong?

ganeshie8 (ganeshie8):

looks ok to me, did u get A = 80pi ?

OpenStudy (anonymous):

i feel stupid. i was putting in a negative area. can't have a negative area....oops.

ganeshie8 (ganeshie8):

:)